n/2-3n/4+5n/6=21




Answer by stanbon(75887)

n/2-3n/4+5n/6 = 21
-----
Multiply thru by 12 to get rid of the fractions.
-------------------
6n - 9n + 10n = 12*21
---------------------
7n = 12*21
----
n = 12*3
-----
n = 36
-------------
Cheers,
Stan H.
------------